Circular No. 6499

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 Q1 (TABUR) This comet began to fade rapidly around Oct. 20, and has become in-

creasingly less condensed as it approaches perihelion on Nov. 3. T. Yusa, Kogota, Japan, obtained a 100-s CCD exposure with a 0.20-m reflector on Oct. 29.41 UT that shows a diffuse, uncondensed, elongated image of total mag ∼ 9. J. E. Bortle, Stormville, NY, reports that a visual observation on Oct. 29.99 showed an elongated “streak” aligned in p.a. 20^{o}–200^{o} (with the comet’s head located toward the northeast end of the streak but showing very little condensation). H. Mikuž, Ljubljana, Slovenia, reports that CCD photometry with a 0.20-m f/2 Baker-Schmidt camera on Oct. 30.74 shows a coma of magnitude V = 8.5 and size ∼ 13′ × 5′ aligned along p.a. ∼ 10^{o}–190^{o} and a straight tail at least 1^{o}.2 long on 5-min V and R exposures; the inner coma is almost uniform without central condensation. W. Offutt, Cloudcroft, NM, notes that the “head” is broad and poorly defined with no apparent condensation on four 50-s CCD frames (0.60-m reflector) obtained under poor conditions during Oct. 31.06–31.07. Total visual magnitude estimates: Oct. 13.75, 5.2 (M. Plsek, Lelekovice, Czech Republic, 7×50 binoculars); 19.16, 5.6 (C. E. Spratt, Victoria, BC, 11×80 binoculars); 20.23, 6.6 (A. Pereira, Cabo da Roca, Portugal, 9×34 binoculars); 22.17, 7.6 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 23.19, 7.8 (G. Glitscher, Darmstadt, Germany, 20×80 binoculars); 29.76, 8.1 (R. J. Bouma, Groningen, The Netherlands, 15×80 binoculars); 29.99, 9.4 (Bortle, 0.32-m reflector); 30.80, 9.3 (Pereira, 0.25-m reflector).

SUPERNOVA 1996bo IN NGC 673 I. Sato, National Astronomical Observatory of Japan, reports the fol-

lowing magnitude and precise position measured from a CCD image obtained by H. Fukushima, N. Yamamoto, and himself at Mitaka on Oct. 30.49 UT: I = 16.3, α = 1^{h}48^{m}22^{s}.862, δ = +11^{o}31′15′′.13 (equinox J2000.0; correction of systematic error from GSC to PPM system by Δα = +0′′.31 ± 0′′.35 and Δδ = −0′′.79 ± 0′′.25), which is 6′′.14 east and 1′′.58 south of the galaxy’s center.

Corrigendum. On IAUC 6497, line 24, for Patat et al. 1996, 278, read Patat et al. 1996, MNRAS 278,

1996 October 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green